tv sets: according to the Nielsen company, the mean number of TV sets in a U.S. household in 2013 was 2.24. Assume the standard deviation is 1.4. a sample of 90 households is drawn. Find the 30th percentile of the sample mean.
Answer #1

= 2.24

= 1.4

n = 90

SE = /

= 1.4/ = 0.1476

30th Percentile corresponds to area = 0.50 - 0.30 = 0.20 from mid value to Z on LHS.

Table of Area Under Standard Normal Curve gives Z = - 0.525

So,

Z = - 0.525 = ( - 2.24)/0.1476

So,

=2.24 - (0.525 X 0.1476) = 2.1625.

So,

Answer is:

2.1625